DIFFERENTIAL GEOMETRY/PDE SEMINAR

Monday, August 11, 2003 Padelford C-36 11:00 am (Note: Special Day and Time!)

The conformally invariant operators on differential forms and applications

Rod Gover (University of Auckland)

I will sketch an explicit geometric construction which gives a large class of conformally invariant differential operators acting between form density bundles. We anticipate that the construction is complete in that there are no other operators between form density bundles. In the case of “true forms” (i.e. density weight zero) these operators have remarkable properties. They lead to a new family of conformally invariant elliptic complexes and are linked in particular to the Q-curvature.

This is joint work with Tom Branson.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.